OpenStudy (yousef1231238):

Jean and Mark are going to fill a pool with 2 different sized hoses. Jean can fill the pool in 8 hours, while Mark can complete it in 12 hours. Their supervisor thinks that the job will take 10 hours to complete if they work together. Explain each step in solving this equation, and determine if the supervisor is correct or not.

OpenStudy (evoker):

Jean can fill 1/8th of the pool in an hour while mark can fill 1/12 th of the pool in an hour, that should help you get started.

OpenStudy (yousef1231238):

1/8+1/12?

OpenStudy (evoker):

that will give you the fraction they can fill in an hour yes

OpenStudy (yousef1231238):

so what do I do after?

OpenStudy (evoker):

well if that was 1/2 for instance it would take them 2 hours, so basically invert that answer and you will get the time, the general form is \[\frac{ 1 }{ a }+\frac{ 1 }{ b }=\frac{ 1 }{ c }\]

OpenStudy (yousef1231238):

so 3/24+2/24=5/24

OpenStudy (evoker):

Yep then flip it for 24/5 or 4.8 hours.